# Verifying LLM Inference to Detect Model Weight Exfiltration

### v2

@@ 1 Introduction
As Large Language Models (LLMs) become increasingly valuable assets, the need for robust defenses to protect them grows in parallel. Inference servers are of particular concern as they host the model weights themselves, and a single compromised server can cause severe damage – either through the exfiltration of proprietary weights or by returning undesirable or manipulated outputs during inference.
Malicious adversaries can exploit many potential channels, from infiltrating development environments and stealing backups to abusing side channels. While conventional security mechanisms can mitigate some of these risks, defending the inference pipeline poses a unique challenge. To support normal operation, inference providers must allow vast volumes of data—often hundreds of gigabytes per day—to flow in and out of their data centers [nevo2024securing]. This open communication channel creates an unavoidable tension between accessibility and security.
For attackers, as the amount of user data increases, so does the attractiveness of exfiltrating information through user data through steganography—the practice of concealing information within seemingly innocuous data, such as embedding hidden payloads in generated text. In this work, we propose a scheme to monitor inference traffic to detect potential exfiltration through steganography. Figure 1 depicts two contrasting scenarios: normal inference and compromised inference with potential exfiltration.
@@
Our work rests on a simple but powerful observation: LLM inference traffic follows a well-defined generative process, and any steganographic attempt to encode hidden information must, by necessity, deviate from expected model behavior. The primary challenge is that LLMs naturally exhibit non-determinism—even identical runs of the same model and code can yield slightly different outputs due to stochastic sampling and floating-point variance. In this work, we provide methods for quantifying the likelihood of different tokens under this valid non-determinism, as a way to potential exfiltration.  [^1] It’s worth noting that our scheme protects against a adversary with access to a malicious inference stack, but does not provide defense against attacks where a model reveals its own weights as intended behavior of the model.
In this work, we introduce two methods for estimating the likelihood that each generated token was honestly produced by the claimed model (Token-DiFRand Token-IPT-DiFR). This enables us to measure deviations from expected model behavior and derive information-theoretic bounds on the amount of information that even a capable adversary could exfiltrate through normal inference traffic. We propose a detection scheme that can be integrated into larger model security systems.
@@
Beyond technical design, we offer a framework for reasoning about model-weight exfiltration from both security-engineering and economic perspectives. Intuitively, exfiltration attempts that move large amounts of data quickly are easier to detect and respond to, whereas extremely slow leaks may be economically irrelevant to either party. We highlight three key operational regimes:
@@
    Setting 1: Fast exfiltration: Extremely rapid leaks that occur before human intervention is possible—for example, if an adversary compromises many endpoints and distributes shards of the model simultaneously (e.g., 10,000 servers each leaking 1/10,000th of the weights).
@@
    Setting 2: Intermediate exfiltration: Leaks over days or weeks, allowing incident-response teams to detect and intervene in time.
@@
    Setting 3: Slow exfiltration: Gradual leakage over months or years (e.g., 1% of weights per month), which may render the stolen model economically obsolete by the time it is fully reconstructed.
@@
In this work, our goal is to make this “undetectable” regime as narrow and costly as possible. In practice, many datacenter operators have response times on the order of hours to days. We demonstrate that our proposed techniques can extend effective response windows from days to months or years, providing at least a $100\times$ reduction in feasible exfiltration rates.
@@
This work makes the following contributions:
@@
    Formalizes model-weight exfiltration as a security game and derives strong bounds on exfiltration capacity when using the systems we propose. We formalize the adversary-warden interaction in Experiment 3.2 and prove three bounds under varying assumptions: a general bound for adversaries with seed access (Theorem 4.1) showing at most $r$ bits/token leakage when the threshold is set to $r$, plus tighter bounds when adversaries lack seed access (Theorems 4.2 and 4.3).
@@
    We propose the Fixed-Seed Sampling Likelihood (FSSL) as a method for estimating likelihood of being sampled, and we introduce two estimates for this, Token-DiFR (FSSL-GM), and Token-IPT-DiFR (FSSL-IPT), whic can be used for LLMs that sample with the Gumbel-Max trick and Inverse Probability Transforms, respectively.
@@
    Proposes a decoupled verification server architecture that separates security from performance, enabling high-assurance computation on isolated hardware/software stacks without constraining inference throughput. The architecture requires minimal trust assumptions (detailed in Appendix D) and has dual-use value for both security monitoring and reliability/bug detection.
@@
    Validates empirically on production systems We evaluate exfiltration rates on Llama-3.1-8B, Llama-3.2-3B, Qwen1.5-MoE-7B, Qwen3-30B-A3B. For Qwen3-30B-A3B we demonstrate a $0.5\%$ exfiltrable information rate with $10^{-2}\%$ FPR, a ${\geq}200\times$ slowdown in attack. For a 1 TB model with 10001000Gb of traffic per day, this extends exfiltration time from ${\sim}1$ days to >200{>}200 days.
@@
2 Preliminaries
@@ 2.1 Basic Setting (LLM inference)
Transformer inference consists of two phases: prefill processes the input prompt in a single parallel forward pass, while decode autoregressively generates output tokens one at a time. Though both phases have similar asymptotic complexity, prefill achieves significantly higher throughput in practice (typically 3–10×) because it is compute-bound and highly parallelizable, whereas decode is memory-bandwidth bound due to repeatedly loading the key–value cache [erdil2025inferenceeconomicslanguagemodels]. This makes prefill particularly advantageous for verification or evaluation tasks that do not require autoregressive generation.
@@ 2.2 Sampling From Large Language Models
Sampling from a Large Language model can be thought of in two steps: 1. generating the probability distribution, and then 2. sampling a token from it. For step 1, Given logits $z\in\mathbb{R}^{V}$ over a vocabulary of size $V$, the model induces a categorical distribution $\pi_{i}=\textrm{softmax}(z)_{i}=\exp(z_{i})/\sum_{j=1}^{V}\exp(z_{j})$.. To sample from a probability distribution, many modern LLMs sample tokens using either the Inverse Probability Transform (e.g., Ollama [ollama_docs], HuggingFace [huggingface]) or the Gumbel-Max Trick (e.g., VLLM [vllm]). After applying temperature scaling and filtering (e.g., top-$k$, top-$p$, or min-$p$), the model converts logits into probabilities via the softmax function.
@@
Given a fixed random seed $\sigma$, a uniform random number $u\sim\mathrm{Uniform}(0,1)$ is drawn for each token. The algorithm then computes the cumulative sum of the token probabilities $\vec{p}$ and selects the first token whose cumulative mass exceeds $u$. Conceptually, this corresponds to inverting the cumulative distribution function (CDF) of $\vec{p}$. The method is exact, efficient, and deterministic when both $\vec{p}$ and $\sigma$ are fixed—repeated sampling under identical conditions will always yield the same token.
@@
1:Seed $\sigma$, probability vector $\mathbf{p}\in[0,1]^{V}$ with $\sum_{t=1}^{V}p_{t}=1$
@@
4:return $\min\{t\in\{1,\dots,V\}:C_{t}>u\}\triangleright$ binary search: find smallest $t$ s.t. $C_{t}>u$
@@
The Gumbel-Max Trick provides an equivalent but reparameterized way to sample from $\vec{p}$. Given a fixed random seed $\sigma$, for each token $i$, an independent Gumbel noise term $g_{i}\sim\mathrm{Gumbel}(0,1)$, determined by $\sigma,$ is added to the (log) probability:
$z_{i}=\log p_{i}+g_{i}.$
The selected token is the index of the maximum perturbed score, $\arg\max_{i}z_{i}$. Intuitively, this converts sampling from a categorical distribution into a deterministic argmax over noisy logits. Like the inverse transform, it is reproducible under a fixed seed $\sigma$, since the Gumbel variables $g_{i}$ are deterministically derived from $\sigma$. This trick is particularly convenient for differentiable relaxations (e.g., Gumbel-Softmax) and for implementations optimized for vectorized or GPU-based sampling.
@@
1:Seed $\sigma$, probability vector $\vec{p}\in[0,1]^{V}$ (over tokens $1..V$, $\sum_{i}p_{i}=1$)
@@
4:return $t$

### v3

@@ 1 Introduction
As Large Language Models (LLMs) become increasingly valuable assets, the need for robust defenses to protect them grows in parallel. Inference servers are of particular concern as they host the model weights themselves, and a single compromised server can cause severe damage: either through the exfiltration of proprietary weights or by returning undesirable or manipulated outputs during inference.
Malicious adversaries can exploit many potential channels, from infiltrating development environments and stealing backups to abusing side channels. While conventional security mechanisms can mitigate some of these risks, defending the inference pipeline poses a unique challenge. To support normal operation, inference providers must allow vast volumes of data to flow in and out of their data centers, often hundreds of gigabytes per day [NLK+24]. This open communication channel creates an unavoidable tension between accessibility and security.
For attackers, as the amount of user data increases, so does the attractiveness of exfiltrating information through user data through steganography, the practice of concealing information within seemingly innocuous data, such as embedding hidden payloads in generated text. In this work, we propose a scheme to monitor inference traffic to detect potential exfiltration through steganography. Figure 1 depicts two contrasting scenarios: normal inference and compromised inference with exfiltration.
@@
Our work rests on a simple but powerful observation: LLM inference traffic follows a well-defined generative process, and any steganographic attempt to encode hidden information must, by necessity, deviate from expected model behavior. The primary challenge is that LLMs naturally exhibit non-determinism; even identical runs of the same model and code can yield different outputs due to stochastic token prediction and floating-point variance. In this work, we provide methods for quantifying the likelihood of different tokens under this valid non-determinism, as a way to prevent model exfiltration. [^1] It’s worth noting that our scheme protects against an adversary with access to a malicious inference stack, but does not provide defense against attacks where a model reveals its own weights as intended behavior of the model.
In this work, we introduce two methods for estimating the likelihood that each generated token was honestly produced by the claimed model (Token-DiFR and Token-IPT-DiFR), which are concurrently developed by [KRR+25]. This enables us to measure deviations from expected model behavior and derive information-theoretic bounds on the amount of information that even a capable adversary could exfiltrate through normal inference traffic. We propose a detection scheme that can be integrated into larger model security systems.
@@
Beyond technical design, we offer a framework for reasoning about model-weight exfiltration from both security-engineering and economic perspectives. Intuitively, exfiltration attempts that move large amounts of data quickly are easier to detect and respond to, whereas extremely slow leaks may be economically irrelevant to either party. We highlight three regimes:
@@
    Setting 1: Fast exfiltration: An extremely rapid leak that occur before human intervention is possible. For example, if an adversary compromises many endpoints and exfiltrates data from all of them simultaneously (e.g., 10,000 servers each leaking a 1/10,000th fraction of the weights).
@@
    Setting 2: Intermediate exfiltration: A leak occurs over days or weeks, allowing incident-response teams to detect and potentially intervene.
@@
    Setting 3: Slow exfiltration: Gradual leakage over months or years, which may render the stolen model no longer valuable by the time it is fully reconstructed.
@@
This work focuses on this “undetectable” regime and designs detection tools that provably limit its effectiveness. Based on our estimates, our proposed techniques can extend effective response windows from days to months or years, providing at least a $100\times$ reduction in feasible exfiltration rates.
@@
Inference verification is the task of confirming that a given text output was produced by a specific model running correct inference, rather than by a tampered, substituted, or otherwise divergent process. In this work, we show that inference verification can be applied to the specific threat of steganographic weight exfiltration. We introduce a specific method of inference verification and show that it can be used to bound the capacity available to an adversarial provider trying to encode information in model outputs. This work makes the following contributions:
@@
    Formalizes model-weight exfiltration as a security game and derives strong bounds on exfiltration capacity when using the systems we propose. We formalize the exfiltration-detection interaction in Experiment 3 and prove three bounds under varying adversarial models (Theorems 5.1, 5.2, and 5.3). Along with these adversarial models, we propose architecture which can realize and limit potential adversaries in practice.
@@
    We propose the Fixed-Seed Sampling Likelihood (FSSL) as a method for estimating likelihood of being sampled, and we introduce two estimates for this, Token-DiFR (FSSL-GM), and Token-IPT-DiFR (FSSL-IPT), which can be used for LLMs that sample with the Gumbel-Max trick and Inverse Probability Transforms, respectively.
@@
    Proposes a decoupled verification server architecture that separates security from performance, enabling high-assurance computation on isolated hardware/software stacks without constraining inference throughput. The architecture requires minimal trust assumptions (detailed in Appendix E) and has dual-use value for both security monitoring and reliability/bug detection.
@@
    Validates empirically on production systems We evaluate exfiltration rates on Llama-3.1-8B, Llama-3.2-3B, Qwen1.5-MoE-7B, Qwen3-30B-A3B. For Qwen3-30B-A3B we demonstrate a $0.5\%$ exfiltrable information rate with $10^{-2}\%$ FPR, a ${\geq}200\times$ slowdown in attack. For a 1 TB model with 10001000Gb of traffic per day, this extends exfiltration time from ${\sim}1$ day to >200{>}200 days.
@@
2 Related Works
There is a lot of work in the space of detecting data exfiltration; in this work we specifically focus on approaches that hide information within a channel containing user data, as this is a novel attack surface area within LLM inference. This ties directly into existing literature on steganography and watermarking, as well as basic forms of content filtering.
2.1 Steganography and Watermarking
This section provides a brief overview of ideas and techniques in steganography and watermarking, which are related to the capabilities of the exfiltration adversary in our threat model.
Steganography
Steganography is the problem of communicating covertly over a channel by embedding hidden information into otherwise innocuous messages. A steganographic scheme consists of an encoding algorithm that maps a secret message and (optionally) a key to a cover object, and a decoding algorithm that recovers the hidden message from the resulting output.
As a concrete example, one may embed a secret message into the least significant bits of an image’s pixels. The modified image appears visually unchanged, yet a recipient who knows the embedding rule can recover the hidden message. More generally, modern steganographic schemes are analyzed with respect to an adversary who attempts to distinguish genuine channel outputs from stego outputs.
In this work, our goal is to detect steganographic schemes used to hide model-weight exfiltration attempts. We leverage standard definitions and results from the steganography literature when understanding the limitations of our proposal. For brevity, we defer formal syntax and foundational results to Appendix B. For other cryptographic primitives, such as pseudorandom functions and hash functions, we refer the reader to Katz and Lindell [KL14].
Watermarking.
Watermarking is a related technique often used to encode a single bit of information within a covertext: $1$ if the watermark is present and 0 if the watermark is not. The embedded information can serve many purposes, such as asserting ownership, tracing provenance, or verifying authenticity.
Conceptually, steganography and watermarking lie on a spectrum, and it may be useful to interpret steganography as repeatedly running a watermarking scheme to encode more than a single bit. In this light, while our work focuses on the task of steganography (and its detection), the connection to watermark allows us to pull from theoretical bounds in the watermarking literature.
There is a recent but long line of work in language model watermarking [AARve, KGW+23b, CGZ24, FGJ+25, KTH+23, CG24, GM25, CHS25]. However, most works operate in a model that require the watermarker and detector to share only a small key with the goal of making the watermark as robust as possible. This means that even the highest throughput schemes only embed about $1$ bit per “block,” where a block is text that is being watermarked [^2] The length of a “block” depends on the scheme and security parameter. In practice, we can roughly think of this as requiring 128-bits of empirical entropy. [CHS25].
In contrast, language model steganography [KJG+21] has a higher throughput rate, since it assumes text is unchanged. These schemes can almost reach capacity, embedding as many bits as the entropy of the response.
2.2 Model Weight Exfiltration
In the context of discussing model-weight exfiltration, a recent framework [NLK+24] introduces the notions of security levels from 1 to 5, which increase in defensive capacity. SL1 is a system that can likely thwart amateur attempts. SL2 is a system that can likely thwart most professional opportunistic efforts by attackers that execute moderate-effort or nontargeted attacks. SL3 is a system that can likely thwart cybercrime syndicates or insider threats. SL4 is system that can likely thwart most standard operations by leading cyber-capable institutions. SL5 is a system that could plausibly be claimed to thwart most top-priority operations by the top cyber-capable institutions.
To set the grounding for this work, we will focus primarily on creating a scheme for moving from SL3 to SL4 and building the basis for SL5 security, by increasing the time required to to covertly exfiltrate from minutes and hours to days and years.
2.3 Inference Verification
As introduced above, inference verification is the task of confirming that text outputs were produced by a specific model running correct inference. Existing approaches span a wide range of trust assumptions and computational costs, from cryptographic guarantees to statistical tests.
Cryptographic methods. Zero-knowledge proofs (ZKPs) offer the strongest guarantees, allowing a verifier to confirm that a computation was performed correctly without re-executing it. Recent work has applied ZKPs to neural network inference [CZL+25, SLZ24], but the computational overhead remains prohibitive: proof generation for LLaMA-2-scale models requires on the order of hundreds to thousands of seconds per token, making these approaches impractical for production deployments.
Internal state fingerprinting. A more tractable class of methods verifies inference by comparing intermediate activations between a provider’s run and a verifier’s replay. LOGIC [SVS+25] collects top-$k$ log probabilities at sampled positions, while TOPLOC [OFP+25] captures top-$k$ activation indices and values from the final hidden layer, encoding them as a polynomial to reduce communication costs. Both methods work by replay: the verifier re-executes inference on the provider’s output sequence and checks that the resulting activations match. However, internal state fingerprinting only confirms that a sequence is consistent with a given forward pass, not that it was actually produced by the model’s sampling procedure. A malicious provider could generate arbitrary text, post-hoc compute matching fingerprints, and pass verification, making these methods proofs of reconstructability rather than generation authenticity.
Distributional tests. [GLG25] frame inference verification as a two-sample test between a trusted distribution and an API distribution, using Maximum Mean Discrepancy with string kernels. [ZYQ+25] propose a rank-based uniformity test that checks whether target API outputs are statistically consistent with a reference model’s distribution over many samples. These approaches are fully black-box but verify only distributional similarity, not per-generation correctness.
Shared seed verification. Concurrently with this work, Karvonen et al. introduce Token-DiFR [KRR+25]. FSSL-GM compares output tokens directly against a trusted reference implementation conditioned on the same seed and is able to detect 4-bit quantization with AUC >0.999>0.999 within 300 output tokens at zero communication overhead. Karvonen et al. also introduce Activation-DiFR which complements this by projecting intermediate activations through random orthogonal projections into compact fingerprints, achieving the same detection accuracy with as few as 2 tokens while reducing communication costs by 25–75% relative to prior methods.
2.4 Existing Model Weight Exfiltration Detection methods
One naive attack to exfiltrate data through user traffic would be to exfiltrate weights directly through output messages, by sending the model weights as plaintext. A first line of defense against such a naive weight exfiltration technique is to detect messages that deviate from ordinary conversational patterns. Two simple, independent approaches can be implemented here.
Syntactic filtering. A basic scheme applies regular expressions and lightweight heuristics to flag messages with suspicious surface patterns, such as long sequences of numbers or other structured encodings. While this approach is trivial to deploy it is also easily evaded by re-encoding information (e.g., mapping numbers to words or distributing bits across multiple tokens).
Perplexity filtering. A slightly more sophisticated defense uses a reference LLM trained on benign message traffic to compute sequence-level perplexity, a standard measure of how well a model predicts a sequence. Perplexity is a function of the log-likelihood of each token, where lower values indicate that the model finds the text highly predictable, while higher values suggest the text is surprising or out-of-distribution. Messages with unusually high perplexity are flagged as potential exfiltration attempts [KAR24]. This can detect outputs that are syntactically normal but statistically implausible under the benign distribution, though adaptive adversaries can still craft messages that maintain natural perplexity while encoding information.
Both schemes are simple to run and provide a low-cost first defense, but they leave a substantial gap between benign and adversarial outputs. A powerful adversary can easily create a steganographic technique that only modifies words and tokens that leave the perplexity of the whole sequence low.
3 LLM Inference Preliminaries
Because our exfiltration attacks operate by modifying the inference procedure, we begin by reviewing how modern LLMs generate text. In standard implementations, inference is designed to be stochastic: given a fixed model and a fixed random seed, sampling-based decoding should be reproducible.
At the algorithmic level, this suggests a clean separation between deterministic computation (forward passes, logit computation) and controlled randomness (token sampling that takes a seed). However, in practice, hardware and software effects introduce additional sources of nondeterminism; and as a result, even with a fixed seed, repeated runs of inference often do not produce identical outputs.
In this section, we formalize which components of inference are intended to be deterministic, which are intentionally randomized, and which remain nondeterministic despite the use of a seed.
@@ 3.1 Basic Setting (LLM inference)
Transformer inference consists of two phases: prefill processes the input prompt in a single parallel forward pass, while decode autoregressively generates output tokens one at a time. Though both phases have similar asymptotic complexity, prefill achieves significantly higher throughput in practice (typically 3–10×) because it is compute-bound and highly parallelizable, whereas decode is memory-bandwidth bound due to repeatedly loading the key–value cache [ERD25]. This makes prefill particularly advantageous for verification or evaluation tasks that do not require autoregressive generation.
@@ 3.2 Sampling From Large Language Models
Sampling from a Large Language Model can be thought of in two steps: (1) generating a probability distribution, and (2) sampling a token from it. For step (1), given logits $z\in\mathbb{R}^{|\mathcal{T}|}$ over a vocabulary $\mathcal{T},$ and a temperature parameter $T$, the model induces a categorical distribution
$p_{i}^{(T)}=\mathrm{softmax}\!\left(\frac{z}{T}\right)_{i}=\frac{\exp(z_{i}/T)}{\sum_{j=1}^{|\mathcal{T}|}\exp(z_{j}/T)}.$
The temperature $T$ controls the sharpness of the distribution: as $T\to 0$, the distribution concentrates on the maximum-logit token (approaching greedy decoding), while larger $T$ produces a flatter distribution and increases randomness. After optional filtering steps such as top-$k$, top-$p$, or min-$p$, the model samples from the resulting categorical distribution. To sample a token from a probability distribution, many modern LLMs use either the Inverse Probability Transform (e.g., Ollama [32], HuggingFace [WDS+20]) or the Gumbel-Max Trick (e.g., vLLM [KLZ+23]). This sampling process relies on a pseudorandom generator, which takes as an input a sampling seed $\sigma$ and counter $i$ as an input, and returns a value $x$.
@@
From a fixed random seed $\sigma$, a (pseudo)random number $u\sim\mathrm{Uniform}(0,1)$ is drawn for each token. The algorithm then computes the cumulative sum of the token probabilities $\vec{p}$ and selects the first token whose cumulative mass exceeds $u$. Conceptually, this corresponds to inverting the cumulative distribution function (CDF) of $\vec{p}$. The method is exact, efficient, and deterministic when both $\vec{p}$ and $\sigma$ are fixed; repeated sampling under identical conditions will always yield the same token. Here we use the notation $F(\cdot;r)$ to denote running a randomized function $F$ with randomness $r$, making it (ideally) deterministic. We also use a deterministic hash function $\mathsf{H}$ to expand $\mathsf{seed}$ into reproducible pseudorandomness. [^3] In practice, we may only need a pseudorandom function, since our security theorems only requires that the output is pseudorandom to an adversary.
@@
1:Seed $\mathsf{seed}$, context position $i$, probability vector $\mathbf{p}\in[0,1]^{|\mathcal{T}|}$ with $\sum_{t=1}^{|\mathcal{T}|}p_{t}=1$
@@
4:return $\min\{t\in\{1,\dots,|\mathcal{T}|\}:C_{t}>u\}\triangleright$ binary search: find smallest $t$ s.t. $C_{t}>u$
@@
The Gumbel-Max Trick provides an equivalent but reparameterized way to sample from $\vec{p}$. From a fixed random seed $\sigma$, for the $t$th token, an independent Gumbel noise term $g_{t}\sim\mathrm{Gumbel}(0,1)$, determined by $\mathsf{seed},$ is added to the (log) probability: $z_{t}=\log p_{t}+g_{t}.$ The selected token is the index of the maximum perturbed score, $\arg\max_{t}z_{t}$. Intuitively, this converts sampling from a categorical distribution into a deterministic argmax over noisy logits. Like the inverse transform, it is reproducible under a fixed seed $\mathsf{seed}$, since the Gumbel variables $g_{t}$ are deterministically derived from $\mathsf{seed}$. This trick is particularly convenient for differentiable relaxations (e.g., Gumbel-Softmax) and for implementations optimized for vectorized or GPU-based sampling.
@@
1:Seed $\mathsf{seed}$, context position $i$, probability vector $\vec{p}\in[0,1]^{|\mathcal{T}|}$ (over tokens $\mathcal{T}$, $\sum_{t}p_{t}=1$)
@@
4:return $\hat{t}$
Sampling Multiple Tokens.
We use the notation $\mathcal{D}^{\mathsf{H}}_{\theta,\mathcal{H}}(\mathsf{seed})$ to denote the honest sampling of a full response for a specific model $\theta,$ context $\mathcal{H},$ seed $\mathsf{seed},$ and hash function (or PRF) $\mathsf{H}.$ This algorithm first samples the distribution $\vec{p}$ using $\theta$ and $\mathcal{H}.$ Next, $\mathcal{D}$ will use one of the two above approaches for selecting a specific token $t$ from the distribution using the randomness from $\mathsf{H}(\mathsf{seed}\|i),$ where $i$ is the current response index. [^4] For one of our bounds, we require that randomness is sampled from $\mathsf{H}(\mathsf{seed}\|i\|\mathcal{H})$, which we call sequential randomness expansion. Then, $t$ is added to the context $\mathcal{H}$ and the process repeats until a termination symbol or response length limit is reached.